A small fan has blades that have a radius of 0.0600 m. When the fan is turned on, the tips of the blades have a tangential acceleration of 19.6 m/s2as the fan comes up to speed. What is the angular acceleration ? of the blades?
Answer #1

tangential acceleration = r* angular acceleration

19.6 = 0.06 * alpha

alpha = 326.67 rad/s^2
Answer #2



angular acc = a/r

= 19.6/0.06

= 326.67 rad/s^2
Answer #3

angular acceleration=tangential acceleration/radius=19.6/0.06=326.667rad per second.